OpenStudy (anonymous):

Point C is located at (1, 2) and point D is located at (−4, −2). Find the y value of the point that is 1 over 4 the distance from point C to point D. 1 0 1.5 2.25

OpenStudy (anonymous):

you want to find the equation to the line first. You can do that with the slope and the y intercept

OpenStudy (anonymous):

the slope (rise/run) will be the y values of C and D (2 and -2) divided by the x values of C and D (1 and -4)

OpenStudy (anonymous):

so it will be 2-(-2) or 4 divided by 1-(-4) or 5 so 4/5 will be your slope

OpenStudy (anonymous):

the slope is the rate of change of x...since the equation is a line, it will be y=4/5x + b (b is the y intercept

OpenStudy (anonymous):

sorry, this is a lot :/

OpenStudy (anonymous):

it ok its really helpful thankyou

OpenStudy (anonymous):

alrighty :)

OpenStudy (anonymous):

plug in one of your points ( C or D) into your almost equation ( y = 4/5 x + b)

OpenStudy (anonymous):

lets say (1,2) so (2) = 4/5*(1) + b and you solve for b

OpenStudy (anonymous):

b will end up being 1 and 1/5 because 2-4/5=1and1/5

OpenStudy (anonymous):

so your equation is y=4/5x+6/5

OpenStudy (anonymous):

I'm sorry @Kali_s this is super long... do you know pythagorean theorum?

OpenStudy (anonymous):

yes is do

OpenStudy (anonymous):

i*

OpenStudy (anonymous):

sweet! so the bottom side will be 5 and the right side will be 4

OpenStudy (anonymous):

4^2 + 5^2 = c^2 16 + 25 = 41 = c^2 so c = Sqrt(41) so you follow that?

OpenStudy (anonymous):

yeaah

OpenStudy (anonymous):

cool :)

OpenStudy (anonymous):

sqrt(41) is the distance between the two points.

OpenStudy (anonymous):

sorry @Kali_s, ignore that last one..man, you're very patient, multiply 1/4 by the whole y distance. or 4

OpenStudy (anonymous):

i like to learn lol and you're helping me out a lot believe me(:

OpenStudy (anonymous):

Because the total rise is 4, so 1/4 of the rise will be just 1 :)

OpenStudy (anonymous):

Since point C is (1,2) you want to add 1 to the y part then plug it in to your equation so plug 3 into the Y of your equation and solve (x,3) or 3 = 4/5x + 6/5

OpenStudy (anonymous):

WOW

OpenStudy (anonymous):

It's asking for the Y value so you already have it :)

OpenStudy (anonymous):

We know that when y=3 it is 1/4 from point C to point D SOO, your answer is 3 :D

OpenStudy (anonymous):

Thanks for your patience @Kali_s!

OpenStudy (anonymous):

welcomee! and thankyou soo much...

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

@Josh55 3 isnt one of my answer choices:(

OpenStudy (anonymous):

oh dear

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

oh, the slope is negative so instead of 2 + 1 being three, it is 2 - 1 so the answer is 1 instead :)

OpenStudy (anonymous):

thankyou soo much

OpenStudy (anonymous):

haha, you're so patient, Thank You :) I will be more patient with people around ME in the future :)

OpenStudy (anonymous):

lol no problem